not really sure what it is called please help
2x/3+15=17

Answer by checkley71(8403)

2x/3+15=17
2x/3=17-15
2x/3=2 cross multiply.
2x=6
x=6/2
x=3 answer.
proof
2*3/3+15=17
2+15=17
17=17